# Suppressing Overestimation in Q-Learning through Adversarial Behaviors

### v1

@@ Abstract
The goal of this paper is to propose a new Q-learning algorithm with a dummy adversarial player, which is called dummy adversarial Q-learning (DAQ), that can effectively regulate the overestimation bias in standard Q-learning. With the dummy player, the learning can be formulated as a two-player zero-sum game. The proposed DAQ unifies several Q-learning variations to control overestimation biases, such as maxmin Q-learning and minmax Q-learning (proposed in this paper) in a single framework. The proposed DAQ is a simple but effective way to suppress the overestimation bias thourgh dummy adversarial behaviors and can be easily applied to off-the-shelf reinforcement learning algorithms to improve the performances. A finite-time convergence of DAQ is analyzed from an integrated perspective by adapting an adversarial Q-learning. The performance of the suggested DAQ is empirically demonstrated under various benchmark environments.
1 Introduction
@@
Q-learning (Watkins & Dayan, 1992) stands as one of the most foundational and widely-adopted reinforcement learning algorithms (Sutton & Barto, 2018) and has been the subject of extensive study in several works (Jaakkola et al., 1994; Tsitsiklis, 1994; Borkar & Meyn, 2000; Zou et al., 2019; Qu & Wierman, 2020; Xiong et al., 2020; Lee, 2023). Although Q-learning is known to converge to the optimal solution, it has been reported in Thrun & Schwartz (1993) that the intermediate iterates of the algorithm may suffer from maximization bias during the learning process, due to the presence of the max operator in its update formulations. Notably, the maximization bias is especially pronounced when the number of actions at a given state is substantial. To mitigate this bias, Double Q-learning was introduced in Hasselt (2010). This approach employs dual estimators and eliminates the max operator in the update, using an effective combination of the two estimators. More recently, the so-called maxmin Q-learning was proposed in Lan et al. (2020) as a means to mitigate overestimation bias. This approach introduces multiple Q-estimators and incorporates the min operator to counterbalance the bias induced by the max operator. A similar concept has also been presented in Fujimoto et al. (2018) to diminish overestimation bias in actor-critic algorithms. This is achieved by employing double Q-estimators and incorporating the min operator between the two estimators to counteract the inherent bias.
@@
The main objective of this paper is to introduce a novel variant of Q-learning – termed Dummy Adversarial Q-learning (DAQ). This variant incorporates an additional dummy adversarial player and offers a unified framework that encompasses existing approaches. Specifically, it includes maxmin Q-learning, as proposed in Lan et al. (2020), and minmax Q-learning, introduced in this paper as a minor variant of the maxmin version, as special cases within this framework. Beyond this, the framework affords a unified perspective, aiding in the comprehension of the existing maxmin and minmax Q-learning algorithms and facilitating the analysis of their finite-time convergence in a unified fashion.
@@
Specifically, DAQ deploys multiple Q-estimators akin to maxmin Q-learning (Lan et al., 2020), yet with additional modifications to the reward function involving the integration of constant shifts, all while preserving the original objective. The core intuition in DAQ stems from the notion that it can be interpreted as introducing a dummy player within the intrinsic Markov decision process, thus establishing a two-player zero-sum Markov game (Shapley, 1953). The dummy player strives to minimize the return without influencing the environment, which subsequently enables effective regulation of overestimation biases in the Q update formulations. Consequently, it has been established that DAQ is equivalent to Minimax Q-learning (Littman, 1994), an algorithm formulated for zero-sum two-player Markov games (Shapley, 1953). Therefore, its convergence can be analyzed by directly applying existing analytical frameworks such as Littman & Szepesvári (1996) and Lee (2023) with minimal alterations, which are also included in this paper.
@@
We anticipate that the principles in DAQ can be integrated into existing value-based reinforcement learning algorithms, potentially enhancing their performance. Finally, we furnish comprehensive empirical experiments to validate the effectiveness of the proposed DAQ across diverse benchmark tasks.
1.1 Related Works
@@
To mitigate overestimation bias, several approaches have been introduced. Double Q-learning, as described in Hasselt (2010), utilizes dual estimators and removes the max operator in the update, employing an effective blend of the two estimators. Maxmin Q-learning presented in Lan et al. (2020), introduces multiple Q-estimators and integrates the min operator to counteract the bias resulting from the max operator. A concept akin to this has been developed in Fujimoto et al. (2018), which involves the use of double Q-estimators and the incorporation of the min operator between the two estimators to neutralize the inherent bias. Bias-corrected Q-learning is proposed in Lee et al. (2013), applying intricate bias-correction terms. Weighted double Q-learning (Zhang et al., 2017) manages a weight between a single estimate and double estimates. Moreover, it was reported in Ren et al. (2021) that Double Q-learning might introduce underestimation biases, potentially leading to multiple non-optimal fixed points.
2 Preliminaries
2.1 Markov Decision Process
@@
For reference, we first briefly introduce the standard Markov decision problem (MDP) (Puterman, 2014), where a decision-making agent sequentially takes actions to maximize cumulative discounted rewards in environments called the Markov decision process. A Markov decision process is a mathematical model of dynamical systems with the state-space ${\cal S}:=\{1,2,\ldots,|{\cal S}|\}$ and action-space ${\cal A}:=\{1,2,\ldots,|{\cal A}|\}$. The decision maker selects an action $a$ with the current state $s$, then the state transits to a state $s^{\prime}$ with probability $P(s,a,s^{\prime})$, and the transition incurs a reward $r(s,a,s^{\prime})$, where $P(s,a,s^{\prime})$ is the state transition probability from the current state $s\in{\cal S}$ to the next state $s^{\prime}\in{\cal S}$ under action $a\in{\cal A}$, and $r(s,a,s^{\prime})$ is the reward function. At time $t\geq 0$, let $s=s_{t},a=a_{t}$ and $s^{\prime}=s_{t+1}$. For convenience, we consider a deterministic reward function and simply write $r(s_{t},a_{t},s_{t+1})=:r_{t},t\in\{0,1,\ldots\}$. A (deterministic) policy $\pi:{\cal S}\to{\cal A}$ maps a state $s\in{\cal S}$ to an action $\pi(s)\in{\cal A}$. The objective of the Markov decision problem (MDP) is to find a (deterministic) optimal policy $\pi^{*}$ such that the cumulative discounted rewards over infinite time horizons is maximized, i.e., $\pi^{*}:=\operatorname*{arg\,max}_{\pi\in\Theta}{\mathbb{E}}\left[\left.\sum_{t=0}^{\infty}{\gamma^{t}r_{t}}\right|\pi\right]$, where $\gamma\in[0,1)$ is the discount factor and $\Theta$ is the set of all admissible deterministic policies. The Q-function under policy $\pi$ is defined as $Q^{\pi}(s,a)={\mathbb{E}}\left[\left.\sum_{t=0}^{\infty}{\gamma^{t}r_{t}}\right|s_{0}=s,a_{0}=a,\pi\right]$, and the optimal Q-function is defined as $Q^{*}(s,a)=Q^{\pi^{*}}(s,a)$ for all $s\in{\cal S},a\in{\cal A}$. Once $Q^{*}$ is known, then an optimal policy can be retrieved by the greedy policy $\pi^{*}(s)=\operatorname*{arg\,max}_{a\in{\cal A}}Q^{*}(s,a)$. Q-learning (Watkins & Dayan, 1992) is one of the most fundamental and popular reinforcement learning algorithms (Sutton & Barto, 2018) to solve MDP, whose Q-estimator update defined by
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[{r(s_{t},a_{t},s_{t+1})+\gamma{{\max}_{a^{\prime}\in{\mathcal{A}}}}Q({s_{t+1}},a^{\prime})-Q(s_{t},a_{t})}\right]$ (1)
@@
where $\alpha>0$ is a step-size. The Q-estimator $Q$ is known to converge to the optimal Q-function $Q^{*}$, by which optimal policy $\pi^{*}$ can be retrieved.
2.2 Two-Player Zero-Sum Markov Game
@@
In a two-player zero-sum Markov game (Shapley, 1953), two agents compete with each other to maximize and minimize a return, respectively. Hereafter, these two agents will be called the user and the adversary. The user’s goal is to maximize the return, while that of the adversary is to hinder the user and minimize the return. We focus on an alternating two-player Markov game, where the user and the adversary take turns selecting their actions. The user takes an action without knowledge of the adversary. After that, the adversary observes the user’s action and then takes its action.
@@
The seminal work by Littman (1994) introduced the following Minimax Q-learning, a Q-learning algorithm designed for zero-sum two-player Markov games:
$Q({s_{t}},{a_{t}},{b_{t}})\leftarrow Q({s_{t}},{a_{t}},{b_{t}})+\alpha\left[{r(s_{t},a_{t},s_{t+1})+\gamma{{\max}_{a\in{\mathcal{A}}}}{{\min}_{b\in{\mathcal{B}}}}Q({s_{t+1}},a,b)-Q({s_{t}},{a_{t}},{b_{t}})}\right],$
where the actions of the user and adversary ${a_{t}},{b_{t}}$ are selected following some behavior policies. Subsequently, Littman & Szepesvári (1996) established the asymptotic convergence of Minimax Q-learning towards the optimal value derived from game theory. Recently, Lee (2023) proposed a finite-time convergence analysis of Minimax Q-learning based on the switching system model developed in Lee et al. (2023). Later in this paper, we provide a new perspective that the proposed Q-learning algorithms can be analyzed based on Minimax Q-learning for solving two-player zero-sum Markov games.
3 Controlling Overestimation Biases
@@
3.1 Maxmin Q-Learning
To control overestimation biases in the naïve Q-learning, maxmin Q-learning was suggested in Lan et al. (2020), which maintains $N$ estimates of the Q-functions, $Q_{i}$, and using the minimum of these estimates for the Q-learning target. The corresponding update is given by
${Q_{i}}({s_{t}},{a_{t}})\leftarrow{Q_{i}}({s_{t}},{a_{t}})+\alpha\left[r(s_{t},a_{t},s_{t+1})+\gamma{\max_{a^{\prime}\in{\mathcal{A}}}}{\min_{j\in\{1,\ldots,N\}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\right]$
for some randomly selected $i$. When $N=1$, the naïve Q-learning is recovered. As $N$ increases, the updates tend to suppress the overestimation biases.
3.2 Double Q-Learning
A double estimator was first introduced in Double Q-learning (Hasselt, 2010) to reduce the overestimation of Q-learning induced by the max operator. The update is given as follows:
${Q_{i}}({s_{t}},{a_{t}})\leftarrow{Q_{i}}({s_{t}},{a_{t}})+\alpha\left[{r(s_{t},a_{t},s_{t+1})+\gamma{Q_{j}}({s_{t+1}},\arg{{\max}_{a^{\prime}\in{\mathcal{A}}}}{Q_{i}}({s_{t+1}},a^{\prime}))-{Q_{i}}({s_{t}},{a_{t}})}\right]$
@@ 3.2 Double Q-Learning
3.3 Twin-Delayed Deep Deterministic Policy Gradient (TD3)
Twin-delayed deep deterministic policy gradient (TD3) (Fujimoto et al., 2018) is an improved actor-critic algorithm, which considered a target that takes the minimum between the two Q-estimates: $Y^{target}=r(s,a,s^{\prime})+\gamma\min_{j\in\{1,2\}}Q_{\theta^{\prime}_{j}}(s^{\prime},\pi_{\phi_{i}}(s^{\prime}))$ for a pair of actors $\pi_{\phi_{i}}$ and critics $Q_{\theta_{i}}$, and targets $\pi_{\phi^{\prime}}$ and $Q_{\theta^{\prime}}$. This modification leads to reduced overestimation biases due to the newly introduced min operator.
4 Proposed Algorithms
4.1 Minmax Q-Learning
First, we introduce a minor modification to maxmin Q-learning, accomplished by simply switching the order of the min and max operators. In particular, consider the following update:
${Q_{i}}({s_{t}},{a_{t}})\leftarrow{Q_{i}}({s_{t}},{a_{t}})+\alpha\left[r(s_{t},a_{t},s_{t+1})+\gamma{\min_{j\in\{1,\ldots,N\}}}{\max_{a^{\prime}\in{\mathcal{A}}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\right]$
for all $i\in\{1,\ldots,N\}$ (synchronous case) or randomly selected $i$ (asynchronous case), where $N$ represents the number of Q-estimators, the selection of which can be flexibly adapted depending on the problem. We call this algorithm minmax Q-learning [^1] Note that our minmax Q-learning is different from Minimax Q-learning (Littman, 1994) which was designed for a two-player Markov game.. Although it represents a subtle modification of maxmin Q-learning, its characteristics have remained unexplored in existing literature until now. In this paper, we provide experimental results demonstrating that the proposed minmax Q-learning can also effectively regulate overestimation biases. Moreover, we will soon demonstrate that minmax Q-learning can be integrated into a broader class of algorithms, named the Dummy Adversarial Q-learning (DAQ), which will be introduced in the following subsection. This integration enables a unified analysis and interpretation framework alongside maxmin Q-learning.
4.2 Dummy Adversarial Q-Learning (DAQ)

### v3

@@ Abstract
The goal of this paper is to propose a new Q-learning algorithm with a dummy adversarial player, which is called dummy adversarial Q-learning (DAQ), that can effectively regulate the overestimation bias in standard Q-learning. With the dummy player, the learning can be formulated as a two-player zero-sum game. The proposed DAQ unifies several Q-learning variations to control overestimation biases, such as maxmin Q-learning and minmax Q-learning (proposed in this paper) in a single framework. The proposed DAQ is a simple but effective way to suppress the overestimation bias through dummy adversarial behaviors and can be easily applied to off-the-shelf value-based reinforcement learning algorithms to improve the performances. A finite-time convergence of DAQ is analyzed from an integrated perspective by adapting an adversarial Q-learning. The performance of the suggested DAQ is empirically demonstrated under various benchmark environments.
Index Terms:
Reinforcement learning, Q-learning, overestimation bias, finite-time convergence analysis, zero-sum game
I Introduction
@@
Reinforcement learning algorithms aim to learn optimal policies by interacting with an environment. Q-learning [20] stands as one of the most foundational and widely-adopted reinforcement learning algorithms [16] and has been the subject of extensive study [4, 18, 1, 24, 13, 23, 6]. Although Q-learning is known to converge to the optimal solution, [17] has reported that the intermediate iterates of the algorithm may suffer from maximization bias during the learning process, due to the presence of the max operator in its update formulations. Notably, the maximization bias is especially pronounced when the number of actions at a given state is substantial. To mitigate this bias, double Q-learning was introduced in [3]. This approach employs dual estimators and eliminates the max operator in the update, using an effective combination of the two estimators. More recently, the maxmin Q-learning [5] was proposed as a method to mitigate overestimation bias. This approach introduces multiple Q-estimators and incorporates the min operator to counterbalance the bias induced by the max operator. A similar concept has also been presented in [2] to diminish overestimation bias in actor-critic algorithms. This is achieved by employing double Q-estimators and incorporating the min operator between the two estimators to counteract the inherent bias.
@@
The main objective of this paper is to introduce a novel variant of Q-learning – termed dummy adversarial Q-learning (DAQ). This variant incorporates an additional dummy adversarial player and offers a unified framework that encompasses existing approaches. Specifically, it includes maxmin Q-learning, as proposed in [5], and minmax Q-learning, introduced in this paper as a minor variant of the maxmin version, as special cases within this framework. Beyond this, the framework affords a unified perspective, aiding in the comprehension of the existing maxmin and minmax Q-learning algorithms and facilitating the analysis of their finite-time convergence in a unified fashion.
@@
DAQ deploys multiple Q-estimators akin to maxmin Q-learning [5], yet with additional modifications to the reward function involving the integration of constant shifts, all while preserving the original objective. The core intuition in DAQ stems from the notion that it can be interpreted as introducing a dummy player within the intrinsic Markov decision process, thus establishing a two-player zero-sum Markov game [15]. The dummy player strives to minimize the return without influencing the environment, which subsequently enables effective regulation of overestimation biases in the Q update formulations. Consequently, it has been established that DAQ is equivalent to minimax Q-learning [9], an algorithm formulated for zero-sum two-player Markov games [15]. Therefore, its convergence can be analyzed by directly applying existing analytical frameworks such as [10] and [6] with minimal alterations, which are also included in this paper.
@@
We anticipate that the principles in DAQ can be integrated into existing value-based reinforcement learning algorithms, potentially enhancing their performance. Finally, we furnish comprehensive empirical experiments to validate the effectiveness of the proposed DAQ across diverse benchmark tasks. Supplementary materials are available in https://arxiv.org/pdf/2310.06286.
I-A Related Works
@@
To mitigate overestimation bias, several approaches have been introduced. Double Q-learning was firstly introduced in the work of [3]. This approach employs two separate Q-estimators, where the max operator is used to select the greedy action from one estimator, while the other estimator is used to evaluate that action, avoiding the maximization bias that would occur if the same estimator was used for both selection and evaluation. Maxmin Q-learning [5] introduces multiple Q-estimators and integrates the min operator to counteract the bias resulting from the max operator. A concept akin to this has been developed in [2], which involves the use of double Q-estimators and the incorporation of the min operator between the two estimators to neutralize the inherent bias. Bias-corrected Q-learning is proposed in [7], applying intricate bias-correction terms. Moreover, it was reported that double Q-learning might introduce underestimation biases, potentially leading to multiple non-optimal fixed points [14].
II Preliminaries
II-A Markov Decision Process
@@
For reference, we first briefly introduce the standard Markov decision problem (MDP) [12], where a decision-making agent sequentially takes actions to maximize cumulative discounted rewards in environments called the Markov decision process. A Markov decision process is a mathematical model of dynamical systems with the state-space ${\mathcal{S}}:=\{1,2,\ldots,|{\mathcal{S}}|\}$ and action-space ${\mathcal{A}}:=\{1,2,\ldots,|{\mathcal{A}}|\}$. The decision maker selects an action $a$ at the current state $s$, then the state transits to a state $s^{\prime}$ with probability $P(s,a,s^{\prime})$, and the transition incurs a reward $r(s,a,s^{\prime})$, where $P(s,a,s^{\prime})$ is the state transition probability from the current state $s\in{\mathcal{S}}$ to the next state $s^{\prime}\in{\mathcal{S}}$ under action $a\in{\mathcal{A}}$. At time $t\geq 0$, let $s=s_{t},a=a_{t}$ and $s^{\prime}=s_{t+1}$. For convenience, we consider a deterministic reward function and simply write $r(s_{t},a_{t},s_{t+1})=:r_{t},t\in\{0,1,\ldots\}$. A (deterministic) policy $\pi:{\mathcal{S}}\to{\mathcal{A}}$ maps a state $s\in{\mathcal{S}}$ to an action $\pi(s)\in{\mathcal{A}}$. The objective of the Markov decision problem (MDP) is to find a (deterministic) optimal policy $\pi^{*}$ such that the cumulative discounted rewards over infinite time horizons is maximized, i.e., $\pi^{*}:=\operatorname*{arg\!max}_{\pi\in\Theta}{\mathbb{E}}\left[\left.\sum_{t=0}^{\infty}{\gamma^{t}r_{t}}\right|\pi\right]$ , where $\gamma\in[0,1)$ is the discount factor and $\Theta$ is the set of all admissible deterministic policies. The Q-function under policy $\pi$ is defined as $Q^{\pi}(s,a)={\mathbb{E}}\left[\left.\sum_{t=0}^{\infty}{\gamma^{t}r_{t}}\right|s_{0}=s,a_{0}=a,\pi\right]$, and the optimal Q-function is defined as $Q^{*}(s,a)=Q^{\pi^{*}}(s,a)$ for all $s\in{\mathcal{S}},a\in{\mathcal{A}}$. Once $Q^{*}$ is known, then an optimal policy can be retrieved by the greedy policy $\pi^{*}(s)=\operatorname*{arg\!max}_{a\in{\mathcal{A}}}Q^{*}(s,a)$. Q-learning [20] is one of the most fundamental and popular reinforcement learning algorithms [16] to solve MDP, whose Q-estimator update defined by
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+\\ \gamma\max_{a^{\prime}\in\mathcal{A}}Q({s_{t+1}},a^{\prime})-Q(s_{t},a_{t})\biggr{]}$ (1)
@@
where $\alpha\in(0,1)$ is a step-size. The Q-estimator $Q$ is known to converge to the optimal Q-function $Q^{*}$, by which optimal policy $\pi^{*}$ can be retrieved.
II-B Two-Player Zero-Sum Markov Game
@@
In a two-player zero-sum Markov game [15], two agents compete with each other to maximize and minimize a return, respectively. Hereafter, these two agents will be called the user and the adversary. The user’s goal is to maximize the return, while that of the adversary is to hinder the user and minimize the return. We focus on an alternating two-player Markov game, where the user and the adversary take turns selecting their actions. The user takes an action without knowledge of the adversary. After that, the adversary observes the user’s action and then takes its action.
@@
The seminal work by [9] introduced the following minimax Q-learning, a Q-learning algorithm designed for zero-sum two-player Markov games:
$Q(s_{t},a_{t},b_{t})\leftarrow Q(s_{t},a_{t},b_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+\\ \gamma\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}}Q(s_{t+1},a,b)-Q(s_{t},a_{t},b_{t})\biggr{]},$
where the actions of the user and adversary $a_{t}$, $b_{t}$ are selected following some behavior policies. Subsequently, [10] established the asymptotic convergence of minimax Q-learning towards the optimal value derived from game theory. Recently, [6] proposed a finite-time convergence analysis of minimax Q-learning based on the switching system model developed in [8]. Later in this paper, we provide a new perspective that the proposed Q-learning algorithms can be analyzed based on minimax Q-learning for solving two-player zero-sum Markov games.
III Controlling Overestimation Biases
@@
III-A Maxmin Q-Learning
To control overestimation biases in the Q-learning, maxmin Q-learning was suggested [5], which maintains $N$ estimates of the Q-functions, $Q_{i}$, and using the minimum of these estimates for the Q-learning target. The corresponding update is given by
$Q_{i}(s_{t},a_{t})\leftarrow Q_{i}(s_{t},a_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+\\ \gamma\max_{a^{\prime}\in\mathcal{A}}\min_{j\in\{1,\ldots,N\}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\biggr{]}$
for some randomly selected $i$. When $N=1$, the Q-learning is recovered. As $N$ increases, the updates tend to suppress the overestimation biases.
III-B Double Q-Learning
A double estimator was first introduced in double Q-learning [3] to reduce the overestimation of Q-learning induced by the max operator. The update is given as follows:
${Q_{i}}(s_{t},a_{t})\leftarrow{Q_{i}}(s_{t},a_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+\\ \gamma{Q_{j}}(s_{t+1},\operatorname*{arg\!max}_{a^{\prime}\in\mathcal{A}}{Q_{i}}(s_{t+1},a^{\prime}))-{Q_{i}}(s_{t},a_{t})\biggr{]}$
@@ III-B Double Q-Learning
III-C Twin-Delayed Deep Deterministic Policy Gradient (TD3)
TD3 [2] is an improved actor-critic algorithm, which considered a target that takes the minimum between the two Q-estimates: $Y^{target}=r(s,a,s^{\prime})+\gamma\min_{j\in\{1,2\}}Q_{\theta^{\prime}_{j}}(s^{\prime},\pi_{\phi_{i}}(s^{\prime}))$ for a pair of actors $\pi_{\phi_{i}}$ and critics $Q_{\theta_{i}}$, and targets $\pi_{\phi^{\prime}}$ and $Q_{\theta^{\prime}}$. This modification leads to reduced overestimation biases due to the newly introduced min operator.
IV Proposed Algorithms
IV-A Minmax Q-Learning
First, we introduce a minor modification to maxmin Q-learning, accomplished by simply switching the order of the min and max operators. In particular, let us consider the following update:
${Q_{i}}(s_{t},a_{t})\leftarrow{Q_{i}}(s_{t},a_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+\\ \gamma{\min_{j\in\{1,\ldots,N\}}}{\max_{a^{\prime}\in\mathcal{A}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\biggr{]}$
for all $i\in\{1,\ldots,N\}$ (synchronous case) or randomly selected $i$ (asynchronous case), where $N$ represents the number of Q-estimators, the selection of which can be flexibly adapted depending on the problem. We call this algorithm minmax Q-learning [^1] Note that our minmax Q-learning is different from minimax Q-learning [9] which was designed for a two-player Markov game.. Although it represents a subtle modification of maxmin Q-learning, its characteristics have remained unexplored in existing literature until now. In this paper, we provide experimental results demonstrating that the proposed minmax Q-learning can also effectively regulate overestimation biases. Moreover, we will soon demonstrate that minmax Q-learning can be integrated into a broader class of algorithms, named the dummy adversarial Q-learning (DAQ), which will be introduced in the following subsection. This integration enables a unified analysis and interpretation framework alongside maxmin Q-learning.
IV-B Dummy Adversarial Q-Learning (DAQ)